### v1

# Exact Hull Reformulation for Quadratically Constrained Generalized Disjunctive Programs

Abstract
Generalized Disjunctive Programming (GDP) is a framework for optimization problems involving discrete decisions and nonlinear constraints. The widely used Hull Reformulation transforms GDPs into Mixed-Integer Nonlinear Programming (MINLP) problems with tighter continuous relaxations. However, it typically employs an $\varepsilon$-approximation of the closure of the perspective function for nonlinear constraints, which can lead to numerical instability and a weaker continuous relaxation.
We present an exact Hull Reformulation for GDPs with quadratic constraints, preserving the original quadratic structure and avoiding relaxation weakening. The method applies to both convex and non-convex constraints and extends earlier approaches developed for second-order cone representable functions. We prove equivalence to the conventional exact Hull Reformulation for quadratic constraints and demonstrate improved computational performance and numerical stability in extensive computational experiments. Benchmarks include random GDP instances, Continuously Stirred Tank Reactor network optimization, $k$-means clustering, and constrained layout optimization problems. The computational results demonstrate improved solution times and numerical stability of the proposed exact Hull Reformulation, making it a preferable approach for GDP problems with quadratic constraints.
Keywords: Generalized Disjunctive Programming, Exact Hull Reformulation, Perspective Function, Quadratic Constraints, Mixed-Integer Nonlinear Programming, Non-convex Generalized Disjunctive Programming.
1 Introduction
Problems involving continuous and discrete variables, including those that require making discrete decisions, can generally be represented in the form of Mixed-Integer Nonlinear Programs (MINLPs). However, formulating problems involving a large number of discrete decisions using traditional MINLP frameworks can become overly complex and counterintuitive. For instance, common tasks in Process Systems Engineering (PSE), such as selecting process units or determining process configurations, often result in cumbersome algebraic constraints and numerous binary variables. When modeled in the classical form of MINLP, such problems can be challenging for practitioners to implement and interpret [1, 2, 3].
To address these limitations, Generalized Disjunctive Programming (GDP) has emerged as a higher-level modeling paradigm that more naturally integrates logical decision-making with algebraic constraints. Introduced by [4], GDP extends Balas’s disjunctive programming framework [5] to include nonlinear constraints and logical relations [6, 1, 2]. In GDP, discrete alternatives, such as choosing among different equipment configurations or operational modes, are explicitly represented through Boolean variables and logical disjunctions (OR conditions) rather than through algebraically encoded binary variables. This approach enables clearer and more intuitive formulations compared to conventional MINLP models, making it much easier to formulate the model and apply the framework to practical challenges in real-world scenarios [7].
A GDP problem can be formally expressed as follows [4, 6]:
$\begin{aligned} \min_{\mathbf{x},\mathbf{Y}}\quad&f(\mathbf{x})\\ \text{s.t.}\quad&g(\mathbf{x})\leq 0,\\ &\bigvee_{i\in D_{k}}\begin{bmatrix}Y_{ik}\\ h_{ik}(\mathbf{x})\leq 0\end{bmatrix},\quad\forall\,k\in K,\\ &\underset{i\in D_{k}}{\underline{\bigvee}}Y_{ik},\quad\forall\,k\in K,\\ &\Omega(\mathbf{Y})=\text{True},\\ &\mathbf{x}^{\ell}\leq\mathbf{x}\leq\mathbf{x}^{u},\\ &\mathbf{x}\in\mathbb{R}^{n},\\ &Y_{ik}\in\{\text{False},\text{True}\},\quad\forall\,k\in K,\;i\in D_{k}.\end{aligned}\qquad\text{(GDP)}$ (1)
In this GDP formulation, $\mathbf{x}\in\mathbb{R}^{n}$ denotes the continuous decision variables and $\mathbf{Y}$ the collection of Boolean variables. The objective function is given by $f:\mathbb{R}^{n}\to\mathbb{R}$. Global constraints are represented by $g:\mathbb{R}^{n}\to\mathbb{R}^{p}$, where $p$ is the number of global inequalities. The condition $g(\mathbf{x})\leq 0$ is understood componentwise, i.e., $g_{j}(\mathbf{x})\leq 0$ for all $j=1,\dots,p$. Each disjunct $i$ within disjunction $k$ is associated with $h_{ik}:\mathbb{R}^{n}\to\mathbb{R}^{q_{ik}}$, where $q_{ik}$ is the number of inequalities in that disjunct. The condition $h_{ik}(\mathbf{x})\leq 0$ is likewise understood componentwise, i.e., $h_{ik,\ell}(\mathbf{x})\leq 0$ for all $\ell=1,\dots,q_{ik}$. The set $K$ indexes all disjunctions, and $D_{k}$ denotes the set of disjuncts within disjunction $k$. The exclusive OR (XOR) condition, $\underset{i\in D_{k}}{\underline{\bigvee}}Y_{ik},$ ensures exactly one Boolean variable $Y_{ik}$ per disjunction $k$ is True. Possible logical relationships among Boolean variables are represented by $\Omega(\mathbf{Y})=\text{True}$. These logical conditions can be systematically translated into algebraic constraints, enabling their integration into conventional optimization frameworks [2].
Although GDPs can be solved directly using specialized logic-based solvers, the more common practice is to reformulate GDPs into MINLP problems and leverage the maturity and broad availability of MINLP solvers [8, 9, 10, 3]. Among the various reformulation strategies, the Hull Reformulation is particularly notable due to its tight continuous relaxation properties, which could enhance solver performance by providing strong lower bounds [11].
One of the major challenges in applying Hull Reformulation to nonlinear constraints in GDP problems arises from numerical instabilities associated with existing methods to reformulate constraints, such as the widely used $\varepsilon$-approximation of the closure of the perspective function [^1] The term $\varepsilon$-approximation of the closure of the perspective function” is often referred to simply as $\varepsilon$-approximation” in this paper. When we discuss properties or drawbacks of $\varepsilon$-approximation”, we are referring to those of the full method (i.e., the application of the $\varepsilon$-approximation of the closure of the perspective function), not only to the fact that it is an approximation rather than an exact function. [12]. Although computationally practical, the $\varepsilon$-approximation complicates the function structure compared to the original constraints in GDP, leading to numerical issues, and, since it is an approximation, it enlarges the feasible region of the continuous relaxation.
Other works have addressed quadratic constraints through second-order cone representations [13], but it can be extended only to convex quadratic constraints. This motivated the present work, which extends this reformulation approach to any quadratic constraint to form an exact Hull Reformulation. The proposed reformulation retains the quadratic functional form of the original constraints and provides an exact description of the Hull Reformulation for disjunctive constraints without relying on approximations. We provide theoretical proofs of equivalence between the proposed exact reformulation, which maintains the functional structure, and the original exact Hull Reformulation for quadratic constraints. We note that the Hull Reformulation is a valid approach to apply to GDPs with both convex and non-convex constraints; thus, we test GDPs with both convex and non-convex quadratic constraints. Extensive computational experiments conducted in this work illustrate that the exact Hull Reformulation outperforms the $\varepsilon$-approximation method in terms of computational efficiency, numerical stability, and solver reliability in various benchmark problems and solvers.
The remainder of the paper is structured as follows. Section 2 reviews main GDP to MINLP reformulation strategies, including the Hull Reformulation and its associated numerical challenges. Section 3 presents the details on quadratically constrained GDPs and their Hull Reformulation, and Section 4 introduces the proposed exact Hull Reformulation approach and proves its equivalence to the standard definition of exact Hull Reformulation. Section 5 analyzes the implications of the proposed reformulation for convex GDPs. Section 6 outlines the comprehensive computational experiments conducted to evaluate reformulation effectiveness across various benchmark problems, including randomly generated instances, a Continuously Stirred Tank Reactor (CSTR) network optimization problem, $k$-means clustering, and a constrained layout problem. Finally, Section 7 concludes the paper with a summary of key findings and suggestions for future research directions.
2 GDP to MINLP Reformulation Strategies
2.1 MINLP Reformulation Basics
GDP problems are commonly solved by reformulating them into MINLP problems. Although several reformulation approaches exist, all typically share a common structure. All MINLP reformulations introduce binary variables to represent Boolean conditions and convert propositional logic into linear constraints, but they differ in how they reformulate the constraints within the disjunctions. The disjunction-independent part of the MINLP reformulation can be presented as follows:
$\begin{aligned} \min_{\mathbf{x},\mathbf{y}}\quad&f(\mathbf{x})\\ \text{s.t.}\quad&g(\mathbf{x})\leq 0,\\ &\sum_{i\in D_{k}}y_{ik}=1\quad\forall\,k\in K,\\ &\mathbf{E}\mathbf{y}\geq\mathbf{e},\\ &\mathbf{x}^{\ell}\leq\mathbf{x}\leq\mathbf{x}^{u},\\ &\mathbf{x}\in\mathbb{R}^{n},\\ &y_{ik}\in\{0,1\}\quad\forall\,k\in K,\,i\in D_{k}.\end{aligned}\qquad\text{(MINLP without Disjunctions)}$ (2)
In this formulation, the Boolean variables $Y_{ik}$ from the original GDP are converted to binary variables $y_{ik}\in\{0,1\}$, which activate or deactivate the corresponding disjunctive constraints. The constraint $\sum_{i\in D_{k}}y_{ik}=1$ represents the exclusive OR (XOR) condition $\underset{i\in D_{k}}{\underline{\bigvee}}Y_{ik}$, ensuring that exactly one disjunct from each disjunction set $D_{k}$ is selected. All logical relationships between disjuncts, originally denoted by $\Omega(\mathbf{Y})=\text{True}$ are translated into linear constraints $\mathbf{E}\mathbf{y}\geq\mathbf{e}$ using established methods [2].
However, the presented MINLP reformulation does not yet specify how the disjunctive constraints $h_{ik}(\mathbf{x})\leq 0$ should be handled. Each MINLP reformulation strategy differs mainly in how it reformulates these constraints. The following subsections address how this challenge is handled by all reformulations used in this paper.
2.2 Big-M Reformulation
In the Big-M reformulation, constraints within each disjunct are activated conditionally using binary variables and large constants (“Big-M”)[14]. Specifically, disjunctive constraints are reformulated as:
$\begin{aligned} h_{ik}(\mathbf{x})\leq M_{ik}(1-y_{ik})\quad\forall\,k\in K,\,i\in D_{k},\\ \end{aligned}\qquad\text{(Big-M Reformulation)}$ (3)
where $M_{ik}$ are sufficiently large constants chosen such that when $y_{ik}=0$ the constraint is always evaluated to values less than $M_{ik}$ in the domain of the problem, and thus the constraint becomes trivially satisfied regardless of the value of $\mathbf{x}$. On the other hand, when $y_{ik}=1$, the constraint reverts exactly to its original form $h_{ik}(\mathbf{x})\leq 0$, activating the corresponding disjunct.
The quality of continuous relaxation in the case of Big-M reformulation strongly depends on the choice of $M_{ik}$. Excessively large values of $M_{ik}$ lead to weaker continuous relaxations and poorer lower bounds, thus increasing computational effort. Conversely, selecting tighter $M_{ik}$ values improves relaxation quality and solver performance but requires careful, problem-specific tuning. Despite yielding potentially weaker relaxation compared to the Hull Reformulation [11], the Big-M reformulation remains an attractive and widely used approach due to its simplicity, ease of implementation, and the fact that it does not increase the number of variables or constraints compared to the original GDP [15].
2.3 Hull Reformulation
The Hull Reformulation offers tighter continuous relaxations and stronger lower bounds for GDP problems compared to other common approaches, such as Big-M. The Hull Reformulation ensures that the continuous relaxation of each disjunction in the resulting MINLP corresponds exactly to the convex hull of all disjuncts in a disjunction, provided the constraints within disjunctions are convex [11, 13, 15].
2.3.1 Reformulation of Disjunctions
The Hull Reformulation of disjunctions can be expressed as follows:
$\begin{aligned} &\mathbf{x}=\sum_{i\in D_{k}}\mathbf{v}_{ik},\quad k\in K\\ &\left(\text{cl}\,\widetilde{h}_{ik}\right)(\mathbf{v}_{ik},y_{ik})\leq 0,\quad k\in K,\,i\in D_{k}\\ &\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}\\ &\mathbf{v}_{ik}\in\mathbb{R}^{n},\quad k\in K,\,i\in D_{k}\\ \end{aligned}\qquad\text{(Hull Reformulation)}$ (4)
The reformulation introduces disaggregated variables $\mathbf{v}_{ik}$ as separate copies of the original variables $\mathbf{x}$, each associated with disjunct $ik$ in disjunction $D_{k}$. Constraints $\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}$ ensure that $\mathbf{v}_{ik}=0$ when the corresponding binary variable $y_{ik}=0$. The functions in disjunctions $h_{ik}(\mathbf{x})$ are replaced by the closure of the perspective function $\left(\text{cl}\,\widetilde{h}_{ik}\right)(\mathbf{v}_{ik},y_{ik})$ [16], which in the domain of Hull Reformulation, restricted to $\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}$, can be expressed in the following way, provided $h(\mathbf{v})$ is defined and finite in the domain of reformulation [17, 11, 18, 12]:
[…]
Clearly, when $y_{ik}=1$, the original constraint $h_{ik}(\mathbf{x})\leq 0$ is recovered precisely. On the other hand, when $y_{ik}=0$, the reformulated constraint simplifies to $0\leq 0$, making the constraint trivially satisfied.
2.3.2 Numerical Challenges and $\varepsilon$-approximation
Equation (4) outlines the general strategy for handling constraints in Hull Reformulation. In the case of linear constraints of the form $A\mathbf{x}\leq b$, instead of using the definition of closure of the perspective function, the reformulation of linear constraints can be expressed simply as $A\mathbf{v}\leq by$ [5], which preserves linearity and exactly analytically describes the convex hull of the disjunction.
However, nonlinear constraints present a numerical challenge since the closure of the perspective function is not defined analytically at $y=0$. To overcome this issue, $\varepsilon$-approximation method (where $\varepsilon$ is a small number, e.g. $10^{-4}$) proposed by [12] is typically employed:
$\left(\text{cl}\,\widetilde{h}\right)(\mathbf{x},y)\approx\left((1-\varepsilon)y+\varepsilon\right)h\left(\frac{\mathbf{x}}{(1-\varepsilon)y+\varepsilon}\right)-\varepsilon\,h(0)(1-y)$ (6)
Although computationally practical, $\varepsilon$-approximation results in exact Hull Reformulation only when $\varepsilon\rightarrow 0$, and however small, any values $\varepsilon>0$ enlarge the feasible region of the continuous relaxation [12], potentially weakening solver effectiveness. Additionally, application of $\varepsilon$-approximation of the closure of the perspective function creates a more complex functional structure compared to the original disjunctive constraints, introducing numerical instabilities, further motivating the development of an exact reformulation approach presented later in this work.
2.3.3 Convex and Non-Convex Constraints
When dealing with convex GDPs, where each disjunctive constraint is convex, the continuous relaxation of Hull Reformulation creates a convex hull of disjuncts in a disjunction [11], which is the tightest possible convex relaxation for the original disjunctive structure. This property is the key theoretical motivation behind the Hull Reformulation and is what makes it particularly attractive.
However, the behavior of the closure of the perspective function at binary values, described above, and thus the validity of reformulation, is independent of convexity, making it a valid reformulation to apply to non-convex GDPs (Appendix A). Applying Hull Reformulation directly to non-convex GDPs results in non-convex MINLP problems that can be solved by global optimization solvers.
Figure 1 illustrates the feasible region of continuous relaxation obtained from applying Hull Reformulation to a non-convex GDP, providing intuitive insight into the relaxation’s structure and its enlargement when using approximations. However, it should be noted that, since in this case the relaxation is not convex, the solver might need to apply its own techniques to obtain convex relaxations during the solution process to compute lower bounds, meaning that the convex region explored by the solver can be larger compared to the non-convex one obtained after applying the Hull Reformulation.
Figure 1: Illustration of the effect of the $\varepsilon$-approximation on the feasible region of the continuous relaxation resulting from the Hull Reformulation applied to a non-convex GDP. The figure compares the exact feasible region with the enlarged region obtained using the $\varepsilon$-approximation, highlighting how the approximation leads to a relaxation that is larger than the exact one and weakens as $\varepsilon$ increases. The Big-M relaxation is not shown, as its feasible region would significantly exceed the plotted area, taking up most of the figure at this scale and axis limits.
2.3.4 Trade-offs and Practical Considerations
The Hull Reformulation significantly improves the quality of the continuous relaxation, leading to tighter lower bounds. However, this comes at the cost of increased model complexity due to the introduction of additional variables and constraints [7]. This trade-off between stronger relaxations and higher dimensionality makes it difficult to determine in general whether the hull reformulation will outperform the Big-M reformulation in terms of solution time, and the results vary on a case-by-case basis [15, 3].
2.4 Binary Multiplication Reformulation
One of the simplest and most straightforward methods to transform the GDP problem into a MINLP formulation is the binary multiplication approach. This method involves directly multiplying each constraint within a disjunct by the binary variable associated with the disjunct, which activates or deactivates the corresponding constraint. Binary multiplication reformulation of disjunctions can be expressed as:
$\begin{aligned} &y_{ik}h_{ik}(\mathbf{x})\leq 0\quad\forall\,k\in K,\,i\in D_{k},\\ \end{aligned}\qquad\text{(Binary Multiplication Reformulation)}$ (7)
In this formulation, when the binary variable $y_{ik}=0$, the corresponding constraint becomes $0\leq 0$, which is trivially satisfied. When $y_{ik}=1$, the original constraint $h_{ik}(\mathbf{x})\leq 0$ is recovered. Therefore, the disjunctive logic of the original GDP formulation is correctly enforced, and the reformulation is valid.
However, the main drawback of this approach is that the continuous relaxation of the reformulation is often non-convex. In fact, the continuous relaxation has the same feasible region as the original MINLP with binary values, when projected to $\mathbf{x}$, providing no additional convex relaxation benefit to the solver. As a result, MINLP solvers must rely on internal convexification techniques to navigate the non-convex feasible space efficiently during the branch-and-bound procedure.
Despite its lack of favorable relaxation properties, the binary multiplication reformulation remains valid and easy to implement. It can be directly applied to GDP problems and passed to a general MINLP solver, which can handle non-convex problems.
3 Quadratically constrained GDP
In this work, we focus on the hull reformulation of GDPs with disjunctive quadratic constraints. Disjunctions of quadratically constrained GDP can be expressed as follows:
$\bigvee_{i\in D_{k}}\begin{bmatrix}Y_{ik}\\ \mathbf{x}^{\top}Q_{jik}\mathbf{x}+\mathbf{c}_{jik}^{\top}\mathbf{x}+d_{jik}\leq 0,\,j\in J_{ik}\end{bmatrix},\quad k\in K$ (8)
where $J_{ik}$ is a set of quadratic constraints in a disjunct, $Q\in\mathbb{R}^{n\times n},\mathbf{c}\in\mathbb{R}^{n},d\in\mathbb{R}$.
Application of Hull Reformulation to the quadratically constrained disjunctions, using $\varepsilon$-approximation of the closure of the perspective function shown in Equation (6), after algebraic simplification yields:
$\left(\operatorname{cl}\,\widetilde{h}\right)(\mathbf{v},y)\approx\frac{\mathbf{v}^{\top}Q_{jik}\mathbf{v}}{(1-\varepsilon)y+\varepsilon}+\mathbf{c}_{jik}^{\top}\mathbf{v}+d_{jik}y\leq 0,\quad j\in J_{ik},\,k\in K,\,i\in D_{k}$ (9)
Thus, it is clear that application of $\varepsilon$-approximation to the quadratically constrained problem results in an MINLP with a more complicated functional structure as the quadratic term is replaced by a fractional term of the form $\frac{\mathbf{v}^{\top}Q_{jik}\mathbf{v}}{(1-\varepsilon)y+\varepsilon}$.
4 Proposed Hull Reformulation
Consider a GDP in which the disjunctions contain quadratic constraints of the form $h(\mathbf{x})\;=\;\mathbf{x}^{\mathsf{T}}Q\mathbf{x}+\mathbf{c}^{\mathsf{T}}\mathbf{x}+d\leq 0,$ with continuous relaxation obtained through the standard exact Hull Reformulation. In lifted space, the feasible set of the continuous relaxation associated with a single constraint is
$S_{1}\;=\;\Bigl{\{}(\mathbf{v},y)\in\mathbb{R}^{n}\times\mathbb{R}\mid y\in[0,1],\;\operatorname{cl}\widetilde{h}(\mathbf{v},y)\leq 0,\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\Bigr{\}},$ (10)
where the closure of the perspective function simplifies to
$\operatorname{cl}\widetilde{h}(\mathbf{v},y)\;=\;\begin{cases}\dfrac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}+\mathbf{c}^{\mathsf{T}}\mathbf{v}+dy,&y>0,\\[6.0pt] 0,&y=0.\end{cases}$ (11)
Multiplying the function $\dfrac{\mathbf{v}^{\mathsf{T}}Q\mathbf{v}}{y}+c^{\mathsf{T}}\mathbf{v}+dy$ by $y$ motivates the alternative set:
$S_{2}$ $=\left\{(\mathbf{v},y)\in\mathbb{R}^{n}\times\mathbb{R}\mid y\in[0,1],\;\mathbf{v}^{\top}Q\mathbf{v}+(\mathbf{c}^{\top}\mathbf{v})y+dy^{2}\leq 0,\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\right\}$ (12)
We demonstrate the equivalence of these two sets in the following proposition:
Proposition 1.
For a quadratic constraint $h(\mathbf{v})\;=\;\mathbf{v}^{\mathsf{T}}Q\mathbf{v}+\mathbf{c}^{\mathsf{T}}\mathbf{v}+d\leq 0,$ define the sets
$S_{1}$ $=\left\{(\mathbf{v},y)\in\mathbb{R}^{n}\times\mathbb{R}\,\middle|\,y\in[0,1],\;\operatorname{cl}\widetilde{h}(\mathbf{v},y)\leq 0,\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\right\}$
where
[…]
and
$S_{2}=\left\{(\mathbf{v},y)\in\mathbb{R}^{n}\times\mathbb{R}\mid y\in[0,1],\;\mathbf{v}^{\top}Q\mathbf{v}+(\mathbf{c}^{\top}\mathbf{v})y+dy^{2}\leq 0,\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\right\}$
Then, the two sets coincide: $S_{1}=S_{2}$
Proof.
Case $y>0$: Since $y$ is positive, the constraint $\frac{\mathbf{v}^{\top}Q\mathbf{v}}{y}+\mathbf{c}^{\top}\mathbf{v}+dy\leq 0$ can be multiplied by $y$ to yield a constraint with the same feasible region:
$\frac{\mathbf{v}^{\top}Q\mathbf{v}}{y}+\mathbf{c}^{\top}\mathbf{v}+dy\leq 0\quad\Longleftrightarrow\quad\mathbf{v}^{\top}Q\mathbf{v}+(\mathbf{c}^{\top}\mathbf{v})y+dy^{2}\leq 0,\quad\text{when }y>0.$
Case $y=0$: The constraints $\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y$ force $\mathbf{v}=\mathbf{0}$ at $y=0$. Substituting $(\mathbf{0},0)$ into the $\mathbf{v}^{\top}Q\mathbf{v}+(\mathbf{c}^{\top}\mathbf{v})y+dy^{2}$ yields 0, identical to $\operatorname{cl}\widetilde{h}(\mathbf{0},0)$. Thus, the feasible points with $y=0$ are the same in both sets.
Since the feasible regions coincide in both subcases, the equivalence $S_{1}=S_{2}$ is established. ∎
Hence, the Hull Reformulation for quadratically constrained GDP problems can be exactly expressed using set $S_{2}$ for each quadratic constraint, avoiding the need for approximations and maintaining quadratic functional structure in the reformulation. Consequently, when applying Hull Reformulation it is proposed to use $S_{2}$ for any disjunctive quadratic constraints. In this case, the exact Hull Reformulation of quadratically constrained disjunctions shown in Equation (8) can be expressed as:
$\mathbf{x}=\sum_{i\in D_{k}}\mathbf{v}_{ik},\quad k\in K$ (13)
$\mathbf{v}_{ik}^{\top}Q_{jik}\mathbf{v}_{ik}+\mathbf{c}^{\top}_{jik}\mathbf{v}_{ik}y_{ik}+d_{jik}y_{ik}^{2}\leq 0,\quad j\in J_{ik},\,k\in K,\,i\in D_{k}$
$\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}$
$\mathbf{v}_{ik}\in\mathbb{R}^{n},\quad k\in K,\,i\in D_{k}$
The proposed exact reformulation provides two major advantages over $\varepsilon$-approximation approaches. First, since it is exact, it yields a tighter continuous relaxation, improving solver performance by delivering stronger lower bounds. Additionally, it preserves the quadratic structure of the problem, thereby enhancing computational efficiency and eliminating numerical instabilities associated with the $\varepsilon$-approximation. This structural preservation enables the use of advancements in Mixed-Integer Quadratically Constrained Programming (MIQCP) solvers when the original problem is quadratically constrained, which is not the case when $\varepsilon$-approximation is used, which alters the problem structure. Thus, we believe that the exact Hull Reformulation presented herein is preferable for any GDP problems involving quadratic constraints.
The proposed reformulation generalizes earlier results for specific quadratic constraints, such as $y\geq x^{2}$ [18], and for second-order cone–representable quadratic functions [13], applicable to convex quadratic function, and extends the approach to arbitrary quadratic functions, including non-convex ones.
Additionally, we note that this approach can be further generalized to polynomial constraints of arbitrary degree, enabling the construction of a Hull Reformulation in which the resulting expressions retain the same polynomial degree (Appendix B). Alternatively, polynomial constraints can be reformulated into an equivalent problem with quadratic constraints using established techniques [19], to which the proposed exact Hull Reformulation for quadratic constraints can be applied. Thus, using either strategy, the proposed exact Hull Reformulation approach can be extended to polynomial constraints of any degree.
5 Convex Quadratically Constrained GDP
Convex GDP usually refers to a problem in which the objective function and all constraints in the original GDP formulation are convex. Specifically, for quadratically constrained GDPs, convexity implies that the matrices $Q$ associated with quadratic constraints are positive semi-definite ($Q\succeq 0$).
If $Q\succeq 0$ then $S_{1}$ is convex and therefore $S_{2}$ is also convex set by Proposition 1, even though $\mathbf{v}^{\top}Qv+(c^{\top}\mathbf{v})y+d\,y^{2}$ may not be a convex function in $(\mathbf{v},y)$. Some solvers detect convexity via the eigenvalue analysis of each quadratic expression and thus they may misclassify $S_{2}$ as non-convex, depending on the values of $\mathbf{c}$ and $d$.
If for convex quadratically constrained GDP it is desired to obtain a reformulation, that contains only convex functions, easily verified by eigenvalue analysis, it is possible to apply known convexification techniques to do so. Consider the following alternative convex set $S_{3}$, derived from set $S_{2}$ by relaxing the bilinear term $\mathbf{v}y$ through Glover’s linearization [20] and by substitution of $dy^{2}$ with $dy$, which is equivalent at binary $y$, to avoid non-convex function for $d<0$:
$S_{3}=\Big{\{}(\mathbf{v},y,\mathbf{z})$ $\in\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}^{n}\;\Big{|}\;y\in[0,1],\;\mathbf{v}^{\top}Q\mathbf{v}+\mathbf{c}^{\top}\mathbf{z}+dy\leq 0,\;\mathbf{z}\leq\mathbf{x}^{u}y,\;\mathbf{z}\geq\mathbf{x}^{\ell}y,\;$ (14)
$\mathbf{z}\leq\mathbf{v}-\mathbf{x}^{u}(1-y),\;\mathbf{z}\geq\mathbf{v}-\mathbf{x}^{\ell}(1-y),\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\Big{\}}$
The transformed formulation utilizing $S_{3}$ maintains exact equivalence at binary values of $y$, ensures convexity of the functions, and can be verified by convexity detection methods. However, this explicitly convex reformulation may compromise solution efficiency because of the increased number of variables and constraints, as well as a potentially weaker continuous relaxation.
The set $S_{3}$ is included primarily to demonstrate that explicitly convex formulations are achievable and to illustrate one possible method by which solvers might internally manage the non-convex constraint $\mathbf{v}^{\top}Q\mathbf{v}+(\mathbf{c}^{\top}\mathbf{v})y+dy^{2}\leq 0$, when $Q\succeq 0$.
However, this approach can be less efficient because it weakens the continuous relaxation and increases computational overhead. Ideally, convexity detection techniques used by solvers should be applied directly to the original constraints in the GDP formulation, and the knowledge that certain constraints in the reformulation or the entire problem are convex should be retained and passed to a solver. This could eliminate unnecessary convexification steps or redundant relaxations that a solver may perform, thereby improving its computational performance. Unfortunately, most current solvers lack the capability to flag individual constraints or entire models as convex in this way, highlighting an area for improvement in solver design.
We believe that if solvers could recognize or be informed of the convexity of constraints mentioned above, or at least accept information indicating that the entire MINLP is convex when all disjunctive constraints are convex, the computational efficiency of the proposed exact Hull Reformulation for convex GDPs could be additionally improved compared to the results presented in this study.
6 Computational experiments
The primary objective of the computational experiments presented in this paper is to evaluate the performance of the proposed exact Hull Reformulation in comparison to $\varepsilon$-approximation method, which is usually used for any nonlinear constraints. Although the theoretical foundations of the Hull Reformulation have been developed for convex GDPs [16, 11], and most existing literature concentrates on convex GDP formulations![13, 11, 7], as highlighted above, Hull Reformulation is a valid approach to apply to both convex and non-convex problems. Consequently, this study investigates both convex and non-convex GDPs to provide a comprehensive comparison.
The existing literature on non-convex GDPs proposes convexifying the constraints within each disjunct prior to applying techniques such as Hull Reformulation or Big-M to obtain a convex MINLP [21]. However, to the best of our knowledge, no prior work has conducted a comparative analysis between the direct application of the Hull Reformulation and initial constraint convexification. This represents a compelling direction for future research, although it lies beyond the scope of this paper. Similarly, a thorough discussion of the advantages and disadvantages of alternative reformulation methods, such as the Big-M and Binary Multiplication approaches, compared to the Hull Reformulation, especially for non-convex problems, is reserved for future studies. However, the performance results are presented here to provide a comprehensive context.
In this section, we present results of computational experiments for various quadratically constrained GDPs, which were solved using various state-of-the-art general MINLP solvers after applying different reformulation approaches, such as $\varepsilon$-approximation Hull Reformulation, proposed exact Hull Reformulation, Big-M, and Binary Multiplication. All test problems were initially formulated as GDP and reformulated to MINLP using Pyomo (version 6.9.2) [22]. As part of this work, we implemented exact Hull Reformulation for quadratic constraints within Pyomo.GDP [23] by extending its existing Hull Reformulation framework. All computational experiments used the default Pyomo setting of $\varepsilon=10^{-4}$ for $\varepsilon$-approximation approach.
The resulting MINLP formulations were subsequently solved using GAMS (version 49.1.0) [24], employing as sub-solver Gurobi (version 12.0) [25], BARON (version 25.2.1.) [26], and SCIP (version 9.2.0) [27].
All experiments were conducted using a single computational thread on an AMD EPYC 7643 48-Core Processor (96 logical cores), which features 256 MB of L3 cache. The server environment was Red Hat Enterprise Linux 9.5 (Plow), equipped with 1 TB of RAM and 15 GB of swap space.
The following subsections summarize the main findings from the computational experiments across different classes of quadratically constrained GDP problems. These include comparisons between multiple solvers and reformulation strategies. Detailed numerical results and problem generation scripts are available online. [^2] Available at: https://github.com/SECQUOIA/exact_quadratic_hull.
6.1 Random Instances
To evaluate the effectiveness of the proposed Hull Reformulation comprehensively, we conducted computational experiments using randomly generated instances of quadratically constrained GDPs, with quadratic disjunctive constraints, no global constraints, and a quadratic objective function. These problems featured dense matrices and vectors of coefficients, meaning all coefficients were populated with nonzero values. Coefficients were randomly generated from a uniform distribution over the interval [1,1][-1,1]. For the convex problem instances, eigenvalues of the generated random matrices $Q$ were adjusted to be non-negative to ensure positive semidefiniteness of $Q$, thereby guaranteeing convexity of the problem.
Recall that in the formulation presented in Equation (8), $D_{k}$ is a set of disjuncts in a disjunction $k$, $J_{ik}$ is a set of constraints in a disjunct $ik$, $n$ is the number of dimensions of variable $\mathbf{x}$.
The problem generation code allowed for systematic variation in parameters such as the number of dimensions $n$, disjunctions $|K|$, disjuncts per disjunction $|D_{k}|$, constraints per disjunct $|J_{ik}|$, and the number of explicitly guaranteed feasible points. In the experiments carried out, each instance included 10 randomly generated points explicitly ensured to be feasible by adjusting the constant $d_{jik}$ in the constraints.
A total of 240 convex instances were generated with parameters in the following ranges: $3\leq|K|\leq 10,\;10\leq|D_{k}|\leq 15,\;|J_{ik}|=10,\;3\leq n\leq 7$ and 100 non-convex instances with parameters in the following ranges: $3\leq|K|\leq 10,\;10\leq|D_{k}|\leq 15,\;|J_{ik}|=10,\;3\leq n\leq 9$. Generated GDP models were saved to enable execution with different reformulations and solvers.
Figures 2 and 3 show the cumulative number of instances solved within a given wall-clock time for the convex and non-convex random GDP instances, respectively. As illustrated, the proposed exact Hull Reformulation consistently outperforms the $\varepsilon$-approximation approach, solving substantially more instances within the same time.
It is also important to address a significant issue concerning the number of problems solved successfully, particularly with the $\varepsilon$-approximation method. In the conducted computational experiments, the solvers often erroneously identified solutions as optimal within the specified tolerance $(10^{-6})$ to the lower bound or incorrectly labeled instances as infeasible, despite the explicit feasibility of all generated problems. However, for the same problems, other reformulations or alternative solvers identified feasible solutions with significantly better objective values. We attribute these inconsistencies primarily to numerical instabilities associated with the $\varepsilon$-approximation method.
Table 1 summarizes the number of instances in which the solvers returned incorrect results. A solution was considered incorrect if the solver declared the instance infeasible or, without reaching the time limit, reported an optimal solution with objective value relatively larger by more than $10^{-4}$ compared to those obtained by other reformulations, despite the solver’s optimality tolerance being set to $10^{-6}$. The threshold of $10^{-4}$ was chosen to ensure that such cases represent definitively incorrect solutions rather than results only marginally exceeding the specified tolerance. These incorrect solutions were excluded from the performance profile data shown in Figures 2 and 3. This filtering, for example, explains the absence of a line for the $\varepsilon$-approximation in Figure 3(b), since although it solved 81 instances within the time limit, none of the solutions were correct.
Additionally, in the case of SCIP, the solver sometimes reported an error indicating that the problem failed to converge due to numerical instability, rather than providing a result. This is the reason why, in some cases, the total number of reported solutions in Table 1 is less than the total number of tested instances.
Results presented in Table 1 show that the $\varepsilon$-approximation method had the highest frequency of wrong results in tested problems, and the proposed exact Hull Reformulation demonstrated better numerical stability and reported wrong solutions in fewer cases.
In addition to performance plots, Figures 4 and 5 provide direct comparisons of solution times for the same problems solved using the proposed exact Hull Reformulation and the $\varepsilon$-approximation approach. These figures indicate that whenever both methods successfully converged to the same correct solution, the exact Hull Reformulation achieved faster solutions in the vast majority of cases.
In summary, computational experiments on randomly generated quadratically constrained GDPs demonstrate that the proposed exact Hull Reformulation offers not only improved computational efficiency but also enhanced numerical stability, resulting in fewer solver failures and fewer incorrect solutions compared to the $\varepsilon$-approximation method.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 2: Cumulative number of instances solved versus solution time for different reformulations applied to random convex quadratically constrained GDP problems. Results are shown separately for Gurobi, BARON, and SCIP. Each curve represents a distinct reformulation strategy.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 3: Cumulative number of instances solved versus solution time for different reformulations applied to random non-convex quadratically constrained GDP problems. Results are shown separately for Gurobi, BARON, and SCIP. Each curve represents a distinct reformulation strategy.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 4: Solution time for random convex quadratically constrained GDPs using Gurobi, BARON, and SCIP for the same problems, comparing the $\varepsilon$-approximation and the proposed exact Hull reformulation.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 5: Solution time for random non-convex quadratically constrained GDPs using Gurobi, BARON, and SCIP for the same problems, comparing the $\varepsilon$-approximation and the proposed exact Hull reformulation.
Table 1: Solver outcomes across reformulation strategies in convex and non-convex cases
(a) Convex Case
Strategy Optimal Timeout Wrong Total
Solver: Gurobi
BigM 240 0 0 240
Hull ($\varepsilon$-approx.) 102 123 15 240
Hull Exact 240 0 0 240
Binary Mult. 211 29 0 240
Solver: BARON
BigM 239 1 0 240
Hull ($\varepsilon$-approx.) 10 127 103 240
Hull Exact 122 88 30 240
Binary Mult. 198 42 0 240
Solver: SCIP
BigM 240 0 0 240
Hull ($\varepsilon$-approx.) 57 168 7 232
Hull Exact 140 95 0 235
Binary Mult. 112 127 0 239
(b) Non-convex Case
Strategy Optimal Timeout Wrong Total
Solver: Gurobi
BigM 97 3 0 100
Hull ($\varepsilon$-approx.) 47 27 26 100
Hull Exact 81 19 0 100
Binary Mult. 91 9 0 100
Solver: BARON
BigM 100 0 0 100
Hull ($\varepsilon$-approx.) 0 19 81 100
Hull Exact 38 62 0 100
Binary Mult. 98 2 0 100
Solver: SCIP
BigM 94 6 0 100
Hull ($\varepsilon$-approx.) 16 80 0 96
Hull Exact 46 45 0 91
Binary Mult. 52 48 0 100
6.2 Continuously Stirred Tank Reactor network problem
Consider the following superstructure optimization problem involving a network of Continuously Stirred Tank Reactors (CSTRs). The objective is to minimize the total volume of CSTRs in series, in which the autocatalytic reaction $A+B\rightarrow 2B$ takes place. The superstructure allows a recycle stream originating from the last reactor to be returned to any preceding reactor within the series. An illustration of this problem is provided in Figure 6. The details and formulations of the problem were previously described in the literature [28, 9] and are summarized in Appendix C.
Originally, the constraints defining this optimization problem were linear or cubic. Since any polynomial constraint can be systematically represented using quadratic constraints, for the purposes of this work, cubic constraints are reformulated into equivalent quadratic forms, using quadratic equality constraints. Consequently, due to the presence of quadratic equality constraints, the resulting GDP formulation is inherently non-convex. Therefore, this problem allows for testing reformulations in an application, where constraints are linear and non-convex quadratic. This structure makes the problem particularly suitable for evaluating reformulation strategies in applications where constraints are both linear and non-convex quadratic.
Figure 6: Superstructure representation for the continuously stirred tank reactor (CSTR) network problem. The configuration includes $N_{T}$ reactor stages arranged in series, where each stage can either be a reactor or a bypass. The flow proceeds from left to right, and a recycle stream from the last reactor is allowed to return to any upstream stage. This flexible structure enables the optimizer to choose the optimal number of reactors and location of recycle to minimize the total reactor volume. [28, 9].
Computational experiments were conducted to evaluate solution times for the CSTR network optimization problem under various reformulation strategies. In these experiments, the maximum permissible number of reactors in the network was varied, keeping all other model inputs unchanged.
Figure 7 summarizes the computational performance obtained in this work. The clearest difference between reformulations was seen when problems were solved using BARON. As shown in Figure 7(b), the exact Hull Reformulation performed similarly or notably better for all problems in terms of solution time compared to the $\varepsilon$-approximation when solved using BARON. In several instances, the proposed exact formulation provided faster solutions compared to all other methods tested.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 7: Solve time comparison of different reformulations of GDP for the CSTR network problem using Gurobi, BARON, and SCIP.
The results obtained using Gurobi showed a similar trend but with less distinct advantages. Although the exact Hull Reformulation still solved the majority of instances more rapidly compared to $\varepsilon$-approximation, there were cases where the $\varepsilon$-approximation slightly outperformed the exact reformulation. However, the differences in solution times between these reformulations (or in fact between any reformulation tested) were generally minor. These results position the exact Hull Reformulation as a competitive and reliable option for solving this problem with Gurobi, performing better or on par with the $\varepsilon$-approximation while offering theoretical advantages.
Figure 7(c) presents the results obtained using SCIP. For a given reformulation, the solution times fluctuated by orders of magnitude as the number of reactors increased. This behavior indicates high variability in solution time in response to small modifications. Increasing the number of reactors results in a systematic and gradual increase in constraints and variables, while preserving the problem structure, and thus, there should be a corresponding gradual increase in solution time. However, this expected trend was not observed in this case. Therefore, these results do not permit a meaningful comparison between reformulation strategies when using SCIP on this problem.
6.3 $k$-Means Clustering
The $k$-means clustering problem is a classic unsupervised learning task in which $|N|$ data points in $|D|$ dimensions are partitioned into $|K|$ clusters in a way that minimizes the total squared Euclidean distance from each point to its assigned cluster center. Although it is typically solved using heuristics, the problem can also be formulated as a convex GDP with linear and quadratic constraints [29] and solved to global optimality with established reformulation and optimization techniques. This makes it a suitable benchmark problem within the GDP framework [13, 15].
The GDP formulation of the $k$-means clustering problem can be expressed as follows:
$\min_{\mathbf{c},\mathbf{d},Y}$ $\sum_{i\in N}d_{i}$ (15)
s.t. $c_{k-1,1}\leq c_{k,1},\quad k\in\{2,\ldots,|K|\}$
$\bigvee_{k\in K}\left[Y_{ik}\land\left(d_{i}\geq\sum_{j\in D}(p_{ij}-c_{kj})^{2}\right)\right],\quad i\in N$
$\underset{k\in K}{\underline{\bigvee}}Y_{ik},\quad i\in N$
$\mathbf{d}\in\mathbb{R}_{+}^{|N|}$
$\mathbf{c}\in\mathbb{R}^{|K|\times|D|}$
$Y_{ik}\in\{\text{False},\text{True}\},\quad i\in N,k\in K$
Here, $p_{ij}$ is the coordinate of point $i$ in dimension $j$, $c_{kj}$ is the coordinate of cluster center $k$ in dimension $j$, and $d_{i}$ is the squared distance from point $i$ to its assigned center. The constraint $c_{k-1,1}\leq c_{k,1}$ is a symmetry-breaking condition that ensures consistent cluster ordering.
We generated 96 instances of this problem with parameters in the following ranges: $3\leq|K|\leq 5$, $10\leq|N|\leq 17$, and $2\leq|D|\leq 5$. The coordinates of the points were generated randomly by sampling a uniform distribution in the interval [1,1][-1,1].
Figure 8 presents the cumulative number of instances solved as a function of wall‑clock time for each reformulation across the three solvers. In many instances, particularly those solved with SCIP, the time limit was reached, so the profiles are supplemented with cumulative plots showing the absolute optimality gap (the difference between the upper and lower bounds) for these cases. Figure 9 provides a direct comparison of solution times for the same instances across reformulations. A summary of solver outcomes is provided in Table 2.
Across all solvers, the proposed exact Hull Reformulation consistently showed improved numerical stability and outperformed the $\varepsilon$-approximation in terms of the number of instances solved correctly within the time limit. For SCIP and Gurobi, most instances were solved faster with the proposed approach. However, for BARON, the $\varepsilon$-approximation achieved faster solve times for the subset of instances where both approaches converged to the correct solution (Figure 9(b)). However, only 38 out of 96 instances solved using the $\varepsilon$-approximation yielded correct solutions within the time limit, compared to 64 for the proposed reformulation.
We hypothesize that the improved performance of the $\varepsilon$-approximation for this particular problem in BARON may be due to internal convexity detection or solver-specific optimizations tailored to the conventional reformulation and this problem type. Nonetheless, even for BARON from an overall standpoint, the proposed reformulation produced fewer incorrect solutions, and thus more problems solved in the given time limit, affirming its robustness and superior practical value.
These results further demonstrate the advantages of the proposed method for convex GDP formulations involving quadratic constraints such as those arising in $k$-means clustering.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 8: Cumulative number of instances solved versus solution time for different reformulations applied to $k$-means clustering problem. For instances that reached the time limit, cumulative plots are continued with values of absolute optimality gap. Results are shown separately for Gurobi, BARON, and SCIP. Each curve represents a distinct reformulation strategy.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 9: Solution time for $k$-means clustering problem using Gurobi, BARON, and SCIP for the same problems, comparing the $\varepsilon$-approximation and the proposed exact Hull reformulation.
Table 2: Solver outcomes across reformulation strategies for the $k$-means clustering problem
Strategy Optimal Timeout Wrong Total
Solver: Gurobi
BigM 84 12 0 96
Hull ($\varepsilon$-approx.) 34 62 0 96
Hull Exact 82 14 0 96
Binary Mult. 40 56 0 96
Solver: BARON
BigM 91 5 0 96
Hull ($\varepsilon$-approx.) 38 1 57 96
Hull Exact 64 32 0 96
Binary Mult. 74 22 0 96
Solver: SCIP
BigM 19 77 0 96
Hull ($\varepsilon$-approx.) 1 95 0 96
Hull Exact 13 83 0 96
Binary Mult. 17 79 0 96
6.4 Constrained layout problem
The constrained layout involves placing a set of non-overlapping rectangular components within a bounded area to minimize a given objective, the total cost of connections, while satisfying geometric and containment constraints. This problem can be modeled as a convex GDP with both linear and quadratic constraints and is frequently used as a benchmark in GDP literature [30, 12, 13].
In the version of the problem considered here, rectangular 2D components must be positioned within one of several enclosing circular regions, such that the rectangles do not overlap and the total pairwise connection cost is minimized. The objective function employs either an $\ell_{1}$ or $\ell_{2}$ norm to penalize distances between rectangles.
The GDP formulation can be expressed as follows:
$\min_{\delta_{x},\delta_{y},\mathbf{x},\mathbf{y},\mathbf{W},\mathbf{Y}}\quad$ $\sum_{i,j\in N,\,i<j}c_{ij}\cdot\|(\delta_{x_{ij}},\;\delta_{y_{ij}})\|_{p}$
s.t. $\delta_{x_{ij}}\geq x_{i}-x_{j}\quad$ $i,j\in N,\,i<j$
$\delta_{x_{ij}}\geq x_{j}-x_{i}\quad$ $i,j\in N,\,i<j$
$\delta_{y_{ij}}\geq y_{i}-y_{j}\quad$ $i,j\in N,\,i<j$
$\delta_{y_{ij}}\geq y_{j}-y_{i}\quad$ $i,j\in N,\,i<j$
$\left[\begin{array}[]{c}Y_{ij}^{1}\\ x_{i}+L_{i}/2\leq x_{j}-L_{j}/2\end{array}\right]\vee\left[\begin{array}[]{c}Y_{ij}^{2}\\ x_{j}+L_{j}/2\leq x_{i}-L_{i}/2\end{array}\right]$ (20)
$\quad\vee\left[\begin{array}[]{c}Y_{ij}^{3}\\ y_{i}+H_{i}/2\leq y_{j}-H_{j}/2\end{array}\right]\vee\left[\begin{array}[]{c}Y_{ij}^{4}\\ y_{j}+H_{j}/2\leq y_{i}-H_{i}/2\end{array}\right]$ $i,j\in N,\,i<j$ (25)
$\bigvee_{t\in T}\left[\begin{array}[]{c}W_{it}\\ (x_{i}+L_{i}/2-x_{ct})^{2}+(y_{i}+H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\\ (x_{i}+L_{i}/2-x_{ct})^{2}+(y_{i}-H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\\ (x_{i}-L_{i}/2-x_{ct})^{2}+(y_{i}+H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\\ (x_{i}-L_{i}/2-x_{ct})^{2}+(y_{i}-H_{i}/2-y_{ct})^{2}\leq r_{t}^{2}\end{array}\right]$ $i\in N$ (31)
$Y_{ij}^{1}\underline{\vee}Y_{ij}^{2}\underline{\vee}Y_{ij}^{3}\underline{\vee}Y_{ij}^{4}\quad$ $i,j\in N,\,i<j$
$\underset{t\in T}{\underline{\bigvee}}W_{it},\quad$ $i\in N$
$x_{i}^{l}\leq x_{i}\leq x_{i}^{u}\quad$ $i\in N$
$y_{i}^{l}\leq y_{i}\leq y_{i}^{u}\quad$ $i\in N$
$\delta_{x_{ij}},\delta_{y_{ij}}\in\mathbb{R}_{+}\quad$ $i,j\in N,\,i<j$
$x_{i},y_{i}\in\mathbb{R}\quad$ $i\in N$
$Y_{ij}^{1},Y_{ij}^{2},Y_{ij}^{3},Y_{ij}^{4}\in\{\text{False},\text{True}\}\quad$ $i,j\in N,\,i<j$
$W_{it}\in\{\text{False},\text{True}\}\quad$ $i\in N,\,t\in T$
Here, $N$ is the set of rectangles. The variables $x_{i}$ and $y_{i}$ denote the coordinates of the center of rectangle $i$, while $\delta_{x_{ij}}$ and $\delta_{y_{ij}}$ represent the horizontal and vertical distances between rectangles $i$ and $j$. The parameters $L_{i}$ and $H_{i}$ denote the width and height of rectangle $i$, respectively.
The set $T$ contains all enclosing circles, each with center coordinates $(x_{ct},y_{ct})$ and radius $r_{t}$. The binary variables $Y_{ij}^{t}$ encode the relative position of rectangles $i$ and $j$ to enforce non-overlap, while $W_{it}$ indicates whether rectangle $i$ is assigned to circle $t$. The coefficient $c_{ij}$ is the cost of the distance between connected rectangle pairs in the objective.
The objective function minimizes the total pairwise connection cost based on the $\ell_{p}$-norm of the distance between rectangle centers, where $p\in\{1,2\}$. Therefore, depending on the value of $p$, the objective function takes the form:
$\min\sum_{i,j\in N,\,i<j}c_{ij}\cdot\begin{cases}\delta_{x_{ij}}+\delta_{y_{ij}}&\text{if }p=1\\ \sqrt{\delta_{x_{ij}}^{2}+\delta_{y_{ij}}^{2}}&\text{if }p=2\end{cases}$
To assess reformulation performance, we tested six instances of this problem implemented in the Pyomo.GDP example library [^3] https://github.com/Pyomo/pyomo/tree/main/examples/gdp/constrained_layout using both $\ell_{1}$ and $\ell_{2}$ norm. A summary of results for two representative instances (CLay0305 using $\ell_{2}$ norm and CLay0304 using $\ell_{1}$ norm) is provided in Table 3.
For CLay0305, the proposed exact Hull Reformulation solution time was comparable to or greater compared to $\varepsilon$-approximation; however, in the case of SCIP, the $\varepsilon$-approximation reported an incorrect result with large objective and lower bound values. This is attributed to numerical instabilities introduced by the approximation, consistent with observations from other problem types.
In contrast, for CLay0304, the exact Hull Reformulation demonstrated faster solve times with BARON and SCIP compared to the $\varepsilon$-approximation, while in the case of Gurobi, although the solution time was similar, the $\varepsilon$-approximation was faster. With BARON, the $\varepsilon$-approximation hit the 1800s limit with an 11% gap remaining, whereas the exact Hull Reformulation solved the problem in 11.39s to optimality.
Results for all six instances using both $\ell_{1}$ and $\ell_{2}$ norms are provided in Appendix D. While the limited size and number of benchmark problems prevent drawing general conclusions, the observed results suggest that, overall, the exact Hull Reformulation performs on par with the $\varepsilon$-approximation for constrained layout problems.
Table 3: Solver outcomes for CLay0305 and CLay0304 under different norms
(a) CLay0305 with $\ell_{2}$ norm
Strategy Time (s) Objective Lower Bound
Solver: BARON
BigM 53.57 6594.21 6594.21
Hull ($\varepsilon$-approx.) 61.94 6594.21 6594.21
Hull Exact 72.13 6594.21 6594.21
Binary Mult. 78.77 6594.21 6594.21
Solver: Gurobi
BigM 42.85 6594.21 6594.21
Hull ($\varepsilon$-approx.) 218.42 6594.21 6594.21
Hull Exact 858.55 6594.22 6594.21
Binary Mult. 78.53 6594.21 6594.21
Solver: SCIP
BigM 30.50 6594.21 6594.21
Hull ($\varepsilon$-approx.) 21.79 41171.96 41171.96
Hull Exact 20.73 6594.21 6594.21
Binary Mult. 28.77 6594.21 6594.21
(b) CLay0304 with $\ell_{1}$ norm
Strategy Time (s) Objective Lower Bound
Solver: BARON
BigM 0.70 40262.39 40262.39
Hull ($\varepsilon$-approx.) 1808.30 40790.97 36280.70
Hull Exact 11.39 40262.39 40262.39
Binary Mult. 3.84 40262.39 40262.39
Solver: Gurobi
BigM 0.79 40262.39 40262.39
Hull ($\varepsilon$-approx.) 1.79 40262.39 40262.36
Hull Exact 2.13 40262.37 40262.36
Binary Mult. 0.78 40262.39 40262.39
Solver: SCIP
BigM 1.41 40262.39 40262.39
Hull ($\varepsilon$-approx.) 9.59 40262.39 40262.39
Hull Exact 3.28 40262.39 40262.39
Binary Mult. 2.84 40262.39 40262.39
7 Conclusions
This paper proposed a strategy to form an exact Hull Reformulation for any quadratic constraints in GDP problems, effectively addressing the drawbacks associated with the traditional $\varepsilon$-approximation of the closure of the perspective function method. By preserving the quadratic functional structure of constraints and providing an exact reformulation without approximations, the proposed approach improves solver performance, numerical stability, and reliability.
Theoretical results establish that the proposed reformulation is equivalent to the original exact Hull Reformulation, yielding an optimization problem with an identical feasible region. While the conventional exact Hull Reformulation is not expressed in a closed analytical form, our approach provides an analytically defined reformulation that is exact for quadratic functions and therefore directly implementable in practice.
Extensive computational experiments demonstrated that the proposed exact Hull Reformulation approach outperformed the $\varepsilon$-approximation approach in various problem types, including random GDP instances, CSTR network optimization, $k$-means clustering, and constrained layout problems. The results highlighted not only a generally improved solution time, but also a reduction in numerical instabilities and solver failures to provide the correct result.
Furthermore, it is expected that the results obtained could be further improved if the convexity detection methods of the solvers could recognize certain reformulations of the quadratically constrained GDPs as convex, or if the solvers could directly utilize the convexity information obtained before the reformulation. This suggests an area for future development to improve the solvers’ performance.
Several future research directions emerge from this work. First, the performance of the extension of the exact Hull Reformulation to polynomial constraints of higher degree could be thoroughly evaluated, as we did for quadratic constraints in this work. Second, a systematic evaluation could be conducted to compare the direct application of Hull Reformulation to non-convex GDPs against strategies that first convexify the constraints. Third, it would be valuable to compare the performance of the Hull Reformulation with alternative reformulation methods, such as Big-M or Binary Multiplication, when these are applied directly to non-convex GDPs.
Overall, the findings of this study strongly support the adoption of the exact Hull Reformulation as the standard approach for GDP optimization problems involving quadratic constraints.
Acknowledgments
The authors acknowledge and express sincere gratitude to the Davidson School of Chemical Engineering at Purdue University for providing a supportive research environment and funding. The authors thank the GAMS Development Corporation for providing a personal license and access to the BARON solver for our group. We thank Gurobi Optimization for providing access to software through their academic licensing program. We also thank Prof. Can Li for providing access to the hardware.
Appendices
Appendix A From Binary Multiplication to the Hull Reformulation
We outline how the Hull Reformulation for a single disjunction can be systematically derived from the simpler Binary Multiplication reformulation by taking a convex combination of disjuncts. This derivation does not rely on convexity assumptions, demonstrating that the Hull Reformulation is applicable to both convex and non-convex GDPs. It also offers additional insight into the interpretation of the reformulation when applied to non-convex problems. This appendix is adapted from Appendix A of [31].
Consider one disjunction
$\bigvee_{D}\begin{bmatrix}Y_{i}\\ h_{i}(\mathbf{x})\leq 0\end{bmatrix}$
$\underset{i\in D}{\underline{\bigvee}}Y_{i}$
where every $h_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is proper (defined and finite on its domain) and continuously extended at the boundary of the feasible region, and where continuous variables are bounded: $\mathbf{x}^{\ell}\leq\mathbf{x}\leq\mathbf{x}^{u}$.
Binary Multiplication reformulation
The Binary Multiplication reformulation of the disjunction, shown above, activates the constraints by multiplying it with the corresponding binary $y_{i}$:
$y_{i}\,h_{i}(\mathbf{x})$ $\leq 0$ $\forall\;i\in D,$
$\sum_{i\in D}y_{i}$ $=1,$ $y_{i}$ $\in\{0,1\}\quad\forall\;i\in D.$
Because $y_{i}=1$ restores $h_{i}(\mathbf{x})\leq 0$ and $y_{i}=0$ makes the constraint trivially satisfied, the reformulation is exact for binary $y$, and therefore valid. However, its continuous relaxation ($y_{i}\in[0,1]$) provides the same feasible region as the MINLP problem, when projected to $\mathbf{x}$, hence it offers no relaxation advantages and the continuous relaxation problem is equivalent to the original.
Taking a convex combination
Let $\mathbf{x}_{i}$ be a copy of the original variable associated with the feasible region of disjunct $i$. We now form a convex combination of the disjuncts, which yields the convex hull of the disjunction if all $h_{i}(\mathbf{x})$ are convex. Note, however, that convex combinations can also be taken for non-convex sets, but the resulting set will not, in general, be convex, which is illustrated in Figure 1. Thus, these steps remain valid for non-convex problems, although they no longer guarantee the convex hull of the disjunction. The convex combination of disjuncts can then be defined as follows:
$\mathbf{x}=\sum_{i\in D}y_{i}\mathbf{x}_{i},$
$\sum_{i\in D}y_{i}=1,$
$y_{i}\,h_{i}(\mathbf{x}_{i})\leq 0$ $\forall\;i\in D,$
$y_{i}\in[0,1]$
Hull Reformulation
Let $\mathbf{v}_{i}=y_{i}\mathbf{x}_{i}\in\mathbb{R}^{n}$ be a disaggregated copy, scaled by $y_{i}$. Because $\mathbf{x}_{i}$ is bounded by $\mathbf{x}^{\ell}\leq\mathbf{x}_{i}\leq\mathbf{x}^{u}$, $\mathbf{v}_{i}$ is bounded by $\mathbf{x}^{\ell}y_{i}\leq\mathbf{v}_{i}\leq\mathbf{x}^{u}y_{i}$.
The constraint function $y_{i}\,h_{i}(\mathbf{x}_{i})$ that defines the feasible region of each disjunct reduces to zero when $y_{i}=0$, since $h_{i}(\mathbf{x}_{i})$ is assumed to be finite. When $y_{i}>0$, we can apply the change of variables $\mathbf{x}_{i}=\frac{\mathbf{v}_{i}}{y_{i}}$. Also it should be noted that $\mathbf{x}=\sum_{i\in D}y_{i}\mathbf{x}_{i}=\sum_{i\in D}\mathbf{v}_{i}$
Collecting all constraints yields the Hull Reformulation of the disjunction:
$\mathbf{x}=\sum_{i\in D_{k}}\mathbf{v}_{i},$
$\left(\text{cl}\,\widetilde{h}_{i}\right)(\mathbf{v}_{i},y_{i})\leq 0,,i\in D$
$\mathbf{x}^{\ell}y_{i}\leq\mathbf{v}_{i}\leq\mathbf{x}^{u}y_{i}$
$\mathbf{v}_{i}\in\mathbb{R}^{n},i\in D$
where in the domain of reformulation
[…]
Thus, Hull Reformulation can be systematically obtained from Binary Multiplication by taking a convex combination of disjuncts in lifted space. For non-convex $h_{i}(\mathbf{x})$, the construction remains valid, but the continuous relaxation is no longer guaranteed to be convex.
Appendix B Extension of the Exact Hull Reformulation to Polynomial Constraints of Degree $d$
Consider a polynomial disjunctive constraint $h(\mathbf{x})\leq 0$, such that:
$h(\mathbf{x})\;=\;\sum_{k=0}^{d}p_{k}(\mathbf{x}),$
where a scalar polynomial of total degree $d\geq 2$, expressed as a sum of its homogeneous components (parts of the polynomial consisting of all monomials of the same total degree) $p_{k}:\mathbb{R}^{n}\!\to\!\mathbb{R}$ of degree $k$ ($p_{0}$ is the constant term).
The continuous relaxation of the standard Hull Reformulation for this polynomial constraint introduces the following set in the lifted space:
$S_{d}^{(1)}\;=\;\Bigl{\{}(\mathbf{v},y)\in\mathbb{R}^{n}\!\times\![0,1]\;\bigm{|}\;\operatorname{cl}\widetilde{h}(\mathbf{v},y)\leq 0,\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\Bigr{\}},$
where the closure of the perspective function in the domain of the reformulation can be expressed as:
[…]
which, in the case of polynomial constraints, simplifies to:
[…]
Multiplying $\sum_{k=0}^{d}p_{k}(\mathbf{v})\,y^{\,1-k}$ by $y^{\,d-1}$ motivates the alternative set
$S_{d}^{(2)}\;=\;\Bigl{\{}(\mathbf{v},y)\in\mathbb{R}^{n}\!\times\![0,1]\;\bigm{|}\;\sum_{k=0}^{d}p_{k}(\mathbf{v})\,y^{\,d-k}\leq 0,\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\Bigr{\}}.$
We demonstrate the equivalence of these two sets in the following proposition:
Proposition 2.
For any polynomial constraint $h(\mathbf{v})\;=\;\sum_{k=0}^{d}p_{k}(\mathbf{v})\leq 0$ of degree $d\!\geq\!2$, define the sets
$S_{d}^{(1)}\;=\;\Bigl{\{}(\mathbf{v},y)\in\mathbb{R}^{n}\!\times\![0,1]\;\bigm{|}\;\operatorname{cl}\widetilde{h}(\mathbf{v},y)\leq 0,\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\Bigr{\}},$
where
[…]
and
$S_{d}^{(2)}\;=\;\Bigl{\{}(\mathbf{v},y)\in\mathbb{R}^{n}\!\times\![0,1]\;\bigm{|}\;\sum_{k=0}^{d}p_{k}(\mathbf{v})\,y^{\,d-k}\leq 0,\;\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y\Bigr{\}}.$
Then, the two sets coincide: $S_{d}^{(1)}=S_{d}^{(2)}$.
Proof.
Case 1, $y>0$: Because $y^{d-1}>0$, multiplying the defining inequality of $S_{d}^{(1)}$ by $y^{d-1}$ preserves the feasible region and produces the polynomial inequality that defines $S_{d}^{(2)}$. Hence $S_{d}^{(1)}\cap\{y>0\}=S_{d}^{(2)}\cap\{y>0\}$.
Case 2: $y=0$: From the constraints $\mathbf{x}^{\ell}y\leq\mathbf{v}\leq\mathbf{x}^{u}y$ it follows that $y=0$ implies $\mathbf{v}=\mathbf{0}$. Substituting $(\mathbf{v},y)=(\mathbf{0},0)$ into the polynomial expression gives $\sum_{k=0}^{d}p_{k}(\mathbf{0})\,0^{d-k}=0$, matching $\operatorname{cl}\widetilde{h}(\mathbf{0},0)=0$. Therefore, the two sets also coincide on $\{y=0\}$.
Since both cases produce identical feasible regions, the sets are equal. ∎
Therefore, when applying Hull Reformulation, any polynomial disjunctive constraint $\sum_{k=0}^{d}p_{k}(\mathbf{x})\leq 0$, can be reformulated in a way that forms exact hull and preserves polynomial structure as $\sum_{k=0}^{d}p_{k}(\mathbf{v})\,y^{\,d-k}\leq 0$ (sum of monomials of degree $d$), instead of using the closure of the perspective function or its $\varepsilon$‐approximation.
As with the quadratic case ($d=2$), addressed in the paper, the reformulation of polynomials for $d\geq 2$ (i) is exact, guaranteeing a strictly tighter continuous relaxation than $\varepsilon$‐approximation, when $\varepsilon>0$ (ii) retains the original polynomial degree, preserving algebraic structure of the constraints, and (iii) avoids potential numerical instabilities that arise from dividing by small values in $\varepsilon$‐approximations of the closure of the perspective function.
Appendix C Continuously Stirred Tank Reactor (CSTR) Network
This appendix provides a description and the set of equations used to formulate the Continuously Stirred Tank Reactor (CSTR) network optimization problem considered in Section 6, adapted from prior formulations in the literature [28, 9].
The system consists of $N_{T}$ units arranged in series, each of which can be either a reactor or a bypass. The autocatalytic reaction $A+B\rightarrow 2B$ occurs in active reactors.
Each stage $n\in N=\{1,\dots,N_{T}\}$ may operate as a CSTR or be bypassed; exactly one stage receives the fresh feed and exactly one stage receives the recycle stream returned from the outlet of the last stage. Component indices are collected in $I=\{A,B\}$. Variables include molar flow rates $F_{i,n}$ and $FR_{i,n}$, total flow rates $Q_{n}$ and $Q_{FR,n}$, reactor volumes $V_{n}$ (with volume cost $c_{n}$), reaction rates $r_{i,n}$, the recycle flow $(R_{i},Q_{R})$, the product flow $(P_{i},Q_{P})$, and an auxiliary variables $t_{n}$ are used to reduce cubic terms to quadratic form. Boolean variables $YF_{n}$, $YR_{n}$, and $YP_{n}$ denote, respectively, that fresh feed enters stage $n$, recycle enters stage $n$, and stage $n$ operates as a CSTR. All continuous variables are non-negative.
Complete GDP formulation can be presented as follows:
$\min_{F_{i,n},FR,r,t,Q,V,c,YF_{n},YR_{n},YP_{n}}$ $\sum_{n\in N}c_{n}$ (32)
s.t. $F_{i,NT}-F0_{i}-FR_{i,NT}-r_{i,NT}V_{NT}=0$ $\forall\ i\in I$
$Q_{NT}-Q_{F0}-Q_{FR,NT}=0$
$F_{i,n}-F_{i,n+1}-FR_{i,n}-r_{i,n}V_{n}=0$ $\forall\ n\in N\setminus\{N_{T}\},\;i\in I$
$Q_{n}-Q_{n+1}-Q_{FR,n}=0$ $\forall\ n\in N\setminus\{N_{T}\}$
$F_{i,1}-P_{i}-R_{i}=0$ $\forall\ i\in I$
$Q_{1}-Q_{P}-Q_{R}=0$
$P_{i}Q_{1}-F_{i,1}Q_{P}=0$ $\forall\ i\in I$
$0\displaystyle 95\,Q_{P}-P_{B}=$0
$V_{n}-V_{n-1}=0$ $\forall\ n\in N\setminus\{1\}$
$Q_{n}^{2}-t_{n}=0$ $\forall\ n\in N$
$\left[\begin{aligned} &YP_{n}\\ &r_{A,n}t+kF_{A,n}F_{B,n}=0\\ &r_{B,n}+r_{A,n}=0\\ &c_{n}-V_{n}=0\end{aligned}\right]\operatorname*{\underline{\bigvee}}\left[\begin{aligned} &\neg YP_{n}\\ &FR_{i,n}=0\quad(i\in I)\\ &r_{i,n}=0\quad(i\in I)\\ &Q_{FR,n}=0\\ &c_{n}=0\end{aligned}\right]$ $\forall\ n\in N$
$\left[\begin{aligned} &YR_{n}\\ &FR_{i,n}-R_{i}=0\quad(i\in I)\\ &Q_{FR,n}-Q_{R}=0\end{aligned}\right]\operatorname*{\underline{\bigvee}}\left[\begin{aligned} &\neg YR_{n}\\ &FR_{i,n}=0\quad(i\in I)\\ &Q_{FR,n}=0\end{aligned}\right]$ $\forall\ n\in N$
$\underline{\bigvee}_{n\in N}YF_{n},\qquad\underline{\bigvee}_{n\in N}YR_{n}$
$YP_{n}\iff\left(\bigwedge_{j\in\{1,2,..n\}}\neg YF_{j}\right)\vee YF_{n}$ $\forall\ n\in N$
$YR_{n}\Longrightarrow YP_{n}$ $\forall\ n\in N$
$F_{i,n},FR_{i,n},P_{i},R_{i},Q_{n},Q_{FR,n},Q_{F0}\geq 0$ $\forall\ n\in N,i\in I$
$Q_{P},Q_{R},V_{n},c_{n},r_{i,n},t\geq 0$ $\forall\ n\in N,i\in I$
$YF_{n},YR_{n},YP_{n}\in\{\text{False},\text{True}\}$ $\forall\ n\in N$
The first group of equalities enforces component and mass balances (assuming constant density) in the fresh feed stage, interior stages, and the upstream splitter. The purity specification $0.95\,Q_{P}=P_{B}$ guarantees required $0.95mol/l$ concentration of the component $B$ in the outlet. Reactor volumes are constrained to be equal. The auxiliary relation $Q_{n}^{2}=t_{n}$ converts cubic reaction terms to quadratic form. The first disjunction activates the autocatalytic rate law, volume cost, and recycle‐related flows when a stage operates as a reactor. The second disjunction specifies whether or not the recycle stream enters a given stage. Logic conditions enforce that there can be only one unreacted feed and one recycle stream. Also, logic conditions state that a stage and all subsequent stages become a reactor once the fresh feed appears, and that recycle can only return to a reactor stage.
Appendix D Constrained Layout Problem Results
This appendix presents detailed results for the six instances of the constrained layout problem used in the computational experiments discussed in Section 6. These instances originate from the Pyomo.GDP example library and include variations in both objective norms ($\ell_{1}$ and $\ell_{2}$) and layout configurations.
Table 4 summarizes solver outcomes, specifically, the number of optimal solutions, timeouts, and incorrect results, for each of the four reformulation strategies (Big-M, Hull with $\varepsilon$-approximation, exact Hull Reformulation, and Binary Multiplication) across three solvers (BARON, Gurobi, and SCIP). These results provide an overview of the performance and reliability of each reformulation method.
Figures 10 and 11 provide visual comparisons of performance. Figure 10 presents the number of instances solved up to a given wall‑clock. Figure 11 directly compares solution times for each problem instance between the proposed exact Hull Reformulation and the $\varepsilon$-approximation approach.
The proposed exact Hull Reformulation generally matched or outperformed the $\varepsilon$-approximation in terms of solution correctness and numerical stability. In two cases, the $\varepsilon$-approximation produced incorrect solutions, when using SCIP, and in one instance Gurobi reported a wrong solution in the case of the proposed exact Hull Reformulation.
BARON and SCIP showed similar performance in terms of solution time for $\varepsilon$-approximation and exact Hull Reformulation, while in the case of Gurobi $\varepsilon$-approximation showed better results.
Overall, although the limited number and size of the benchmark instances preclude broad generalizations, these results demonstrate that the proposed exact Hull Reformulation is generally competitive with the $\varepsilon$-approximation, and the results may vary depending on the solver.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 10: Cumulative number of constrained layout problem instances solved versus solution time, comparing the $\varepsilon$-approximation and the proposed exact Hull reformulation. Results are shown separately for Gurobi, BARON, and SCIP. Each curve represents a reformulation strategy, highlighting their relative performance in terms of solution speed and number of instances solved.
(a) Gurobi
(b) BARON
(c) SCIP
Figure 11: Solution time comparison for constrained layout problems using Gurobi, BARON, and SCIP, comparing the $\varepsilon$-approximation and the proposed exact Hull reformulation.
Table 4: Solver outcomes across reformulation strategies for the constrained layout problem
Strategy Optimal Timeout Wrong Total
Solver: Gurobi
BigM 12 0 0 12
Hull ($\varepsilon$-approx.) 12 0 0 12
Hull Exact 11 0 1 12
Binary Mult. 12 0 0 12
Solver: BARON
BigM 12 0 0 12
Hull ($\varepsilon$-approx.) 11 1 0 12
Hull Exact 12 0 0 12
Binary Mult. 12 0 0 12
Solver: SCIP
BigM 10 0 0 10
Hull ($\varepsilon$-approx.) 10 0 2 12
Hull Exact 10 1 0 11
Binary Mult. 11 0 0 11